# Building Bridges, Not Walls: Advancing Interpretability by Unifying Feature, Data, and Model Component Attribution

### v2

@@ Abstract
The increasing complexity of AI systems has made understanding their behavior a critical challenge. Numerous methods have been developed to attribute model behavior to three key aspects: input features, training data, and internal model components. However, these attribution methods are studied and applied rather independently, resulting in a fragmented landscape of approaches and terminology. This position paper argues that feature, data, and component attribution methods share fundamental similarities, and bridging them can benefit interpretability research. We conduct a detailed analysis of successful methods of these three attribution aspects and present a unified view to demonstrate that these seemingly distinct methods employ similar approaches, such as perturbations, gradients, and linear approximations, differing primarily in their perspectives rather than core techniques. Our unified perspective enhances understanding of existing attribution methods, identifies shared concepts and challenges, makes this field more accessible to newcomers, and highlights new directions not only for attribution and interpretability but also for broader AI research, including model editing, steering, and regulation.
@@ 1 Introduction
As AI systems grow increasingly complex, understanding their behavior remains a critical challenge (Arrieta et al., 2020; Longo et al., 2024). Researchers have developed methods to explain AI systems by attributing their behavior to three distinct aspects: input features, training data, and internal model components. Feature attribution methods identify influence of input features at test time, revealing which aspects of the input drive the model’s output (Zeiler and Fergus, 2014; Ribeiro et al., 2016; Horel and Giesecke, 2020, 2022; Lundberg and Lee, 2017; Smilkov et al., 2017). Data attribution analyzes how training data shape model behavior during the training phase (Koh and Liang, 2017; Ghorbani and Zou, 2019; Ilyas et al., 2022). Component attribution examines the internal workings of the model by analyzing how specific components, such as neurons or layers in a neural network (NN), affect model behavior (Vig et al., 2020; Meng et al., 2022; Nanda, 2023; Shah et al., 2024). While numerous attribution methods have been developed for each of these three aspects, and some survey papers have been published (Guidotti et al., 2018; Covert et al., 2021; Wang et al., 2024; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024), they have been studied and used rather independently by different communities, creating a fragmented landscape of methods and terminology for similar ideas (Saphra and Wiegreffe, 2024).
Our position is that feature, data, and component attribution methods can be bridged to advance not only interpretability research, by stimulating cross-aspect knowledge transfer, but also broader AI research, including model editing, steering, and regulation. We show that these three types of attribution employ common methods and they differ primarily in perspective rather than core techniques. In the following sections, we first formalize a unified attribution problem that encompasses all three aspects to show these seemingly distinct approaches fall under the same framework (§2). We then examine the evolution of each attribution aspect and analyze its successful methods, revealing how these methods are connected through shared techniques and concepts, including perturbations, gradients, and linear approximations (§3, §4, §5). We summarize the representative attribution methods in these categories in Table 1. Building on this analysis, we present a unified view and illustrate shared concepts (§6.1), identify common challenges (§6.2), and highlight how this unified perspective facilitates cross-aspect knowledge transfer for new research development in interpretability (§6.3), and broader AI research (§6.4). We also discuss alternative views of attribution unification (§7). In summary, we believe that this unified view enhances our understanding of attribution methods, bridges the current fragmented landscape, makes the field more accessible to newcomers, and provides new insights and research directions.
@@ 2 The Attribution Problem
Researchers have developed various attribution methods to analyze model behavior from different perspectives. In this section, we formally introduce three types of attribution problems and show how they fall under a unified framework. Consider a learning problem with -dimensional input features . During training, a dataset of data points: is used to train a model with parameters and components by optimizing the loss function . At test (inference) time, the model generates an output for a new input . For notational simplicity, we omit and “test” and use and when the context is unambiguous. A notation summary is provided in Appendix A. The core objective of all three problems is to attribute the model’s output to different elements and quantify their influence with attribution scores.
Feature attribution quantifies how input features influence model outputs. These features may represent pixels in images, tokens in text, or other domain-specific units. We denote the attribution score of feature as .
Data attribution analyzes how training data shape model behavior. We quantify the influence of each training point through its attribution score .
Component attribution studies the role of model components in generating outputs. The components can have various definitions, such as neurons or layers in a NN. We denote the attribution score of component as .
As illustrated in Figure 1, these three attribution problems share a fundamental connection: they all seek an attribution function that assigns scores to specific elements (features , training points , or components ) for a given test output , differing only in the choice of elements.
| Method | Feature Attribution | Data Attribution | Component Attribution | |
| Perturb | Direct | Occlusions (Zeiler and Fergus, 2014) | LOO (Cook and Weisberg, 1982) | Causal Tracing (Meng et al., 2022) |
@@
| Linear | LIME (Ribeiro et al., 2016) | Datamodels (Ilyas et al., 2022) | COAR (Shah et al., 2024) | |
@@
3 Understanding Feature Attributions
@@
Feature attribution quantifies how individual features of an input influence a model’s output through attribution scores . Applied to model inference at test time, it explains model behavior without altering model parameters. The attribution results can be used to perform feature selection, identify spurious correlations, and justify model predictions to gain user trust. Feature attribution methods can be broadly classified into three categories: perturbation-based methods, gradient-based methods, and linear approximation methods. We discuss some prominent methods in each category below and provide more details in Appendix C.
3.1 Perturbation-Based Feature Attribution
Perturbation-based methods attribute feature importance by measuring how model outputs change when input features are modified and especially removed. They are also referred to as removal-based methods (Covert et al., 2021).
Direct Perturbation represents a straightforward application of perturbation analysis. The pioneering Occlusion method (Zeiler and Fergus, 2014) in computer vision replaces image pixels with grey squares and measures changes in the model’s prediction. The method assumes that occluding crucial pixels will significantly impact the output. For images, pixel attribution scores create a saliency map highlighting the most influential regions. RISE (Petsiuk, 2018) advanced this approach by perturbing multiple image regions and combining their attribution results. The final attribution score weighs each attribution result by the model’s predicted probability for that perturbed image.
@@ 3.1 Perturbation-Based Feature Attribution
Game-Theoretic Perturbation While intuitive, direct perturbation fails to capture synergistic interactions between multiple features. Cooperative game theory addresses this limitation by modeling features as players collaborating toward the model’s output. The Shapley value (Shapley, 1953) provides a foundational solution within this framework and has inspired numerous feature attribution methods (Sundararajan and Najmi, 2020). Computing Shapley value attributions involves measuring a specific type of perturbation: how adding a feature to different feature subsets changes the model’s output compared to the subset alone, known as the marginal contribution of to the subset. The final attribution score captures feature interactions by aggregating these marginal contributions across all possible feature subsets. Although theoretically sound, Shapley value methods face computational challenges as their complexity grows exponentially with feature dimensionality. To overcome this challenge, various approximation methods have been proposed, and Kernel SHAP (or simply SHAP) introduced by Lundberg and Lee (2017) has gained widespread adoption because of its efficient kernel-based approximation.
Perturbation Mask Learning is based on the idea that perturbation of including or excluding features can be viewed as applying a binary mask for each feature. Mask learning methods advance this idea by using learnable masks representing feature inclusion probabilities, which offer more nuanced control compared to binary masks. Dabkowski and Gal (2017) pioneered this approach for image classification by introducing a masking model that generates pixel masks, aiming to identify a minimal set of features that sufficiently maintain the prediction of the original input. The masking model acts as the attribution function , where mask values represent feature attribution scores. While initial training is required, the masking model generates masks through a single forward pass at test time, which significantly improves runtime compared to earlier perturbation methods. For the mask learning methods, the main challenge is balancing feature minimality and predictive power.
3.2 Gradient-Based Feature Attribution
Gradients have emerged as a powerful tool for feature attribution. Gradients of model outputs with respect to input features , , quantify output sensitivity to small input changes (Erhan et al., 2009; Baehrens et al., 2010), measuring feature influence without requiring perturbations. Gradient-based attribution has superior computational efficiency compared to perturbation-based methods. While the latter requires model evaluations for features, gradient-based approaches need only a single or a few forward and backward pass(es) to compute .
@@ 3.2 Gradient-Based Feature Attribution
Gradient-based feature attribution emerged from computer vision, where it gained widespread adoption for generating attribution scores as image saliency maps (Simonyan et al., 2013), also known as sensitivity maps (Smilkov et al., 2017). The “vanilla gradients” method uses the gradients of the output class (log)probability with respect to input pixels as attribution scores (Simonyan et al., 2013). Since then, researchers have proposed numerous enhanced gradient-based methods. For example, Gradients Input (Shrikumar et al., 2017) multiplies gradients with input values, Integrated Gradients (Sundararajan et al., 2017) accumulates gradients along a path from a baseline to the actual input, and Integrated Hessians (Janizek et al., 2021) further extends the analysis to feature interactions by computing the Hessian matrix. These methods leverage different gradient formulations to provide more accurate and stable attribution scores. A notable advancement is SmoothGrad (Smilkov et al., 2017), which generates multiple copies of the input with added Gaussian noise and computes sensitivity maps for each noisy sample. By averaging these maps, SmoothGrad reduces noise while preserving salient features that consistently influence model outputs.
3.3 Linear Approximation for Feature Attribution
Linear approximation methods offer an alternative approach to feature attribution by fitting a simple linear surrogate model around the input of interest. These methods approximate the complex behavior of near a specific input using a linear model, normally in the form of with coefficients and bias . Then the coefficient directly provides a feature attribution score of feature .
@@ 3.3 Linear Approximation for Feature Attribution
LIME (Ribeiro et al., 2016) exemplifies this approach. It samples instances around the input of interest, obtains model predictions for these samples, and fits a sparse linear model to capture the local model behavior. An innovation of LIME is its use of binary indicators (0 or 1) rather than actual feature values as inputs to the linear model, only representing feature inclusion or exclusion. The resulting linear model coefficients directly explain how each feature’s presence influences the approximated model’s output. Later, a variant of LIME called C-LIME improves attribution robustness through its unique neighbor sampling approach for continuous features (Agarwal et al., 2021).
Notably, LIME can also be viewed through a perturbation lens, as it fundamentally perturbs input features to approximate the model’s output. This connection points to a unification: many feature attribution methods can be understood within a common mathematical framework of local function approximation, which we explore next.
3.4 Unifying Feature Attributions via Local Function Approximation
@@
While the original algorithms of feature attribution methods discussed above can be viewed in their respective three categories, many of them can be unified under a common local function approximation framework (Han et al., 2022). Within this framework, a model is approximated around a point of interest in a local neighborhood distribution by an interpretable model using a loss function . Han et al. (2022) show that eight prominent feature attribution methods (Occlusion, KernelSHAP, Vanilla Gradients, Gradients Input, Integrated Gradients, SmoothGrad, LIME, and C-LIME) can be viewed as specific instances of this framework, distinguished only by their unique choices of local neighborhoods and loss functions (Appendix Table 3).
@@
The local function approximation framework (Han et al., 2022) enhances our understanding of feature attribution methods in several important ways. First, it provides conceptual coherence to the field. While different methods appear to have distinct motivations, this framework reveals their shared fundamental goal of local function approximation. Second, placing diverse methods under a single framework enables direct comparisons among them. This comparative lens allows us to better understand their similarities, differences, and behavior, such as why different methods sometimes generate disagreeing or even contradictory explanations for the same model prediction (Krishna* et al., 2024). Third, this unification enables theoretical simplicity. Instead of studying methods seperately, theoretical analyses can be performed using the framework and applied to each method, as shown by the no free lunch theorem and guiding principle in Han et al. (2022). Fourth, the conceptual understanding brought about by unification leads to principled, practical recommendations (Han et al., 2022). Additional details on this unification are provided in Appendix C.6.
4 Understanding Data Attributions
@@
Data attribution studies how the training dataset shapes model behavior. These methods are also known as data valuation, as they help assess the value of data from vendors and content creators. For each training example , an attribution score traces back to the training phase to quantify its influence on the model’s output for a test point . These scores characterize training data properties, help identify mislabeled data, and justify training data values. Like feature attribution, data attribution methods can be organized into three categories: perturbation-based methods, gradient-based methods, and linear approximation methods. We examine prominent methods from each category below, with additional details in Appendix D.
4.1 Perturbation-Based Data Attribution
Perturbation-based data attribution observes the model behavior changes after removing or re-weighting the training data points and subsequently retraining the model, so these methods are also referred to as retraining-based methods (Hammoudeh and Lowd, 2024).
@@ 4.1 Perturbation-Based Data Attribution
Leave-One-Out (LOO) Attribution is a prominent example of this approach, analogous to direct perturbation in feature attribution. The method trains a model on the complete dataset and then separately removes each individual data point and retrains the model. The attribution score for each removed point is determined by the difference in performance between the original and retrained models. The LOO approach has a long history in statistics (Cook and Weisberg, 1982) and has proven valuable for modern AI model data attribution (Jia et al., 2021). It provides valuable counterfactual insights with its main limitation being computational cost, as it requires retraining the model for each data point. Many newer attribution methods can be viewed as efficient approximations of LOO. A natural extension of LOO is to leave a set of data points out to evaluate their collective impact through retraining (Ilyas et al., 2022).
Game-Theoretic Data Attribution represents the successful application of game theory to quantify training sample influence similar to feature attribution. As a direct perturbation method, LOO attribution overlooks interactions between data points, potentially missing subtle influence behaviors (Lin et al., 2022; Jia et al., 2021). Game-theoretic data attribution methods address this by treating training data points as players in a cooperative game, aiming to fairly distribute the model’s performance among training samples. Data Shapley (Ghorbani and Zou, 2019) first applied Shapley values to data attribution by computing each training point’s aggregated marginal contribution across all possible training data subsets. Although theoretically sound, game-theoretic methods face prohibitive computational costs for large datasets, as each marginal contribution requires model retraining and there are possible subsets. Various approximation methods have been proposed to address this challenge, which we discuss in Appendix D.2.
4.2 Gradient-Based Data Attribution
Gradient-based data attribution methods leverage the gradients of the loss with respect to training data and test data to assess the impact of training points on model output . As in Charpiat et al. (2019), simple dot product (GradDot) and cosine similarity (GradCos) between these two gradients are used as similarity measures and consequently attribution scores . Like feature attribution, gradient-based methods often offer greater computational efficiency than perturbation-based methods since they typically require no retraining.
@@ 4.2 Gradient-Based Data Attribution
Influence Function (IF), a classic statistical technique originally developed for analyzing influential points in linear regression (Cook and Weisberg, 1980), has been adapted for modern AI models (Koh and Liang, 2017). IF approximates LOO model parameter changes by Taylor expansion, avoiding explicit retraining. This approximation builds on computing both the gradient and the (inverse) Hessian of the loss with respect to model parameters. IF offers an effective and computationally feasible alternative to LOO, but it also faces several challenges. Its convexity assumptions often do not hold for modern AI models, and its Hessian computation remains expensive for large models. Many methods have been proposed to address these limitations; we discuss IF and these enhancements in detail in Appendix D.3.
Tracing (Training) Path While many gradient-based methods follow IF to compute gradients at the final model parameters, TracIn (Pruthi et al., 2020) introduces a novel approach that traces the influence of training instances throughout the entire training process. The method attributes influence by computing dot products between training and test data gradients at each training step from the initial model parameters to the final model parameters at the end of training, accumulating these to capture a training point’s total influence across the training path. This path tracing approach provides valuable insights into training dynamics while avoiding limitations of LOO and IF, such as assigning identical attribution scores to duplicate training data points. TracIn also offers greater flexibility than IF by eliminating the convexity assumption and Hessian matrix computations. On the other hand, its tracing requires storing intermediate model checkpoints during training, increasing both memory usage and computational costs.
4.3 Linear Approximation for Data Attribution
@@
Datamodel (Ilyas et al., 2022) applies linear approximation to data attribution, similar to LIME in feature attribution. It constructs a linear model with coefficients and vectors as inputs, where each input represents a subset of training data. is learned to map any counterfactual subset of training data to output , where is trained on this subset with the given model architecture and training algorithm. The coefficients of thus represent the attribution scores of the training data points. The method’s counterfactual nature enables evaluation of other attribution methods via the Linear Datamodeling Score (LDS), which compares their attribution score rankings to Datamodel’s ranking. While Datamodel can effectively capture model behavior, constructing this large linear model requires extensive counterfactual data obtained by training model on various subsets, making it computationally intensive. TRAK (Park et al., 2023) addresses these computational challenges by estimating Datamodels in a transformed space where the learning problem becomes convex and can be approximated efficiently. It further improves efficiency through random projection of model parameters and ensemble attribution results of multiple trained models. Though the ensemble approach still requires some model retraining on different subsets, it achieves high estimation accuracy with significantly fewer retraining iterations than Datamodel. Furthermore, both approaches can be viewed as perturbation-based methods, similar to LIME, as they systematically vary training data to construct linear models.
5 Understanding Component Attributions

### v3

# Towards Unified Attribution in Explainable AI, Data-Centric AI, and Mechanistic Interpretability

@@ Abstract
The increasing complexity of AI systems has made understanding their behavior critical. Numerous interpretability methods have been developed to attribute model behavior to three key aspects: input features, training data, and internal model components, which emerged from explainable AI, data-centric AI, and mechanistic interpretability, respectively. However, these attribution methods are studied and applied rather independently, resulting in a fragmented landscape of methods and terminology. This position paper argues that feature, data, and component attribution methods share fundamental similarities, and a unified view of them benefits both interpretability and broader AI research. To this end, we first analyze popular methods for these three types of attributions and present a unified view demonstrating that these seemingly distinct methods employ similar techniques (such as perturbations, gradients, and linear approximations) over different aspects and thus differ primarily in their perspectives rather than techniques. Then, we demonstrate how this unified view enhances understanding of existing attribution methods, highlights shared concepts and evaluation criteria among these methods, and leads to new research directions both in interpretability research, by addressing common challenges and facilitating cross-attribution innovation, and in AI more broadly, with applications in model editing, steering, and regulation.
@@ 1 Introduction
The escalating complexity of AI systems necessitates a deeper understanding of their behavior, giving rise to the field of AI interpretability. A central question in interpretability research is how to attribute an AI system’s behavior to its underlying factors. To address this question, researchers across different domains have developed various attribution methods. Feature attribution (FA), which emerged from explainable AI, quantifies the influence of input features on a model’s output at test time to reveal critical decision drivers (Zeiler and Fergus, 2014; Ribeiro et al., 2016; Horel and Giesecke, 2020, 2022; Lundberg and Lee, 2017; Smilkov et al., 2017). Data attribution (DA), a core approach in data-centric AI, analyzes how training data points shape overall model behavior during the training process (Koh and Liang, 2017; Ghorbani and Zou, 2019; Ilyas et al., 2022). Component attribution (CA), a key branch of mechanistic interpretability, examines how internal model components, such as neurons or layers within a neural network (NN), contribute to model behavior (Vig et al., 2020; Meng et al., 2022; Nanda, 2023; Shah et al., 2024). Although numerous attribution methods and comprehensive survey papers have been published (Guidotti et al., 2018; Covert et al., 2021; Wang et al., 2024; Hammoudeh and Lowd, 2024; Bereska and Gavves, 2024; Rai et al., 2024; Räuker et al., 2023; Vilas et al., 2024), these methods have largely been studied and applied in isolation by their respective communities, resulting in a fragmented landscape of methods and terminology for similar concepts (Saphra and Wiegreffe, 2024).
We take the position that 1) feature, data, and component attribution share core techniques despite their different perspectives, and that 2) this more unified view advances both interpretability, by overcoming common challenges and stimulating cross-attribution innovation, and broader AI research, with applications in model editing, steering, and regulation. In the following sections, we first define the attribution problem and formalize a unified framework that encompasses all three types of attributions (§2). Next, we examine each attribution type, analyzing its popular methods, and demonstrate how all three attribution types share fundamental techniques (i.e., perturbations, gradients, and linear approximations), key concepts, and evaluation criteria (§3). We summarize representative attribution methods based on their techniques in Table 1. Then, we demonstrate how this unified view can advance both interpretability and broader AI research (§4). We also discuss alternative views of the paper’s position (§5). In summary, we believe that this unified view enhances our understanding of attribution methods, bridges the current fragmented landscape, and provides new insights and research directions.
@@ 2 The attribution problem
Various attribution methods have been developed to analyze model behavior from different aspects. In this section, we formally introduce three types of attribution problems and show how they fall under a unified framework. Consider a learning problem with -dimensional input features . During training, a dataset of data points: is used to train a model with parameters and components by optimizing the loss function . At test (inference) time, the model generates an output for a new input . For notational simplicity, we omit and “test” and use and when the context is unambiguous. A notation summary is provided in Appendix A. The core objective of all three problems is to attribute the model’s output to different aspects and quantify their influence with attribution scores.
Feature attribution (FA) quantifies how input features affect the model’s output. A features may be pixels in images, tokens in text, or other domain-specific units with as its attribution score.
Data attribution (DA) analyzes how training data shape model behavior during training. We quantify the influence of each training point through its attribution score .
Component attribution (CA) studies the role of model components in producing model outputs. The component can have various definitions, like a neuron in an NN, with as its attribution score.
As illustrated in Figure 1, these three attribution problems share a fundamental connection: they all seek an attribution function that assigns scores to specific aspects (features , training points , or components ) for a given test output , differing only in the choice of aspects.
| Technique | Technique Details | Feature Attribution | Data Attribution | Component Attribution |
| Perturbation | Direct | Occlusions (Zeiler and Fergus, 2014) | LOO (Cook and Weisberg, 1982) | Causal Tracing (Meng et al., 2022) |
@@
| Linear Approximation | LIME (Ribeiro et al., 2016) | Datamodels (Ilyas et al., 2022) | COAR (Shah et al., 2024) | |
@@
3 Unifying feature, data, and component attribution
3.1 Understanding feature attribution (FA)
@@
FA quantifies how individual features of an input influence the model’s output through attribution scores . Applied to model inference at test time, it explains model behavior without altering model parameters. The attribution results can help with feature selection, model debugging, and improve user trust. FA methods can be broadly classified into three categories: perturbation-based, gradient-based, and linear approximation. We discuss some prominent methods in each category below and provide more details in Appendix C.
@@
Perturbation-Based FA. These methods attribute feature importance by measuring how model outputs change when input features are modified and especially removed (also called removal-based methods (Covert et al., 2021)). Direct perturbation straightforwardly applies this idea, e.g., the pioneering Occlusion method (Zeiler and Fergus, 2014) in computer vision (CV) replaces image pixels with grey squares and measures prediction changes. For images, pixel attribution scores create a saliency map highlighting the most influential regions. RISE (Petsiuk, 2018) advances this method by perturbing multiple image regions and combining the results. While intuitive, direct perturbation fails to capture synergistic interactions between multiple features. Game-theoretic perturbation addresses this limitation by modeling features as game players collaborating toward the model’s output. The Shapley value (Shapley, 1953) provides a foundational solution within this framework and has inspired numerous FA methods (Sundararajan and Najmi, 2020). Computing Shapley value attributions involves measuring a specific type of perturbation: how adding a feature to different feature subsets changes the model’s output compared to the subset alone, known as the marginal contribution of to the subset. The final attribution score captures feature interactions by aggregating these marginal contributions across all possible feature subsets, which can be computationally challenging. Various approximation methods have been proposed for this computation, and KernelSHAP Lundberg and Lee (2017) (or simply SHAP) has gained widespread adoption because of its efficient kernel-based approximation. Since removing input features can be viewed as applying binary masks for features, perturbation mask learning methods use learnable masks representing feature removal probabilities, which offer more nuanced control compared to binary masks. A new mask can be learned for each input (Fong and Vedaldi, 2017) or only once as a global masking model (Dabkowski and Gal, 2017). The masking model acts as the attribution function . While initial training is required, can be faster at test time by generating masks through a single forward pass.
@@
Gradient-based FA. Gradients of model outputs with respect to input features , , quantify output sensitivity to small input changes (Erhan et al., 2009; Baehrens et al., 2010), measuring feature influence without requiring perturbations. Gradient-based methods are more computationally efficient, needing only a single or a few forward and backward pass(es) to compute instead of at least model perturbations for features. Also originating from CV, these methods gained widespread adoption for generating image saliency maps (Simonyan et al., 2013) (also called sensitivity maps (Smilkov et al., 2017)). The “vanilla gradients” method uses the gradients of the output class (log)probability with respect to input pixels as attribution scores (Simonyan et al., 2013). Since then, numerous enhanced gradient-based methods has been proposed. For example, Integrated Gradients (Sundararajan et al., 2017) accumulates gradients along a path from a baseline to the actual input, and Integrated Hessians (Janizek et al., 2021) further extends the analysis to feature interactions by computing the Hessian matrix. SmoothGrad (Smilkov et al., 2017) generates multiple copies of the input with added Gaussian noise and computes sensitivity maps for each noisy sample. It then averages these maps to reduce noise while preserving salient features for model outputs.
@@
Linear approximation for FA. These methods offer an alternative approach to FA by approximating the complex behavior of around an input of interest using a linear model, typically in the form of with coefficients and bias , and using coefficient as the FA score of feature . LIME (Ribeiro et al., 2016) exemplifies this approach. It fits a sparse linear model to capture local model behavior and uses binary indicators (0 or 1) rather than actual feature values as inputs to the linear model, only representing feature exclusion or inclusion. The resulting linear model coefficients explain how each feature’s presence influences the approximated model’s output. Notably, LIME can also be viewed through a perturbation lens, as it fundamentally perturbs input features to approximate the output. This connection suggests that different FA methods can be understood within a common mathematical framework, which we explore next.
3.1.1 Unifying feature attribution methods via local function approximation (LFA)
@@
While FA methods can be grouped in the three categories above, many of them can be unified under the LFA framework (Han et al., 2022). In this framework, a model is approximated around a point of interest in a local neighborhood by an interpretable model using a loss function . Eight prominent FA methods (Occlusion, KernelSHAP, Vanilla Gradients, Gradients Input, Integrated Gradients, SmoothGrad, LIME, and C-LIME) are instances of this framework, differing only in their choices of local neighborhoods and loss functions (Appendix Table 3) Han et al. (2022).
@@
The LFA framework (Han et al., 2022) enhances our understanding of FA methods in several ways. First, it provides conceptual coherence to the field. While different methods appear to have distinct motivations, this framework reveals their shared goal of LFA. Second, placing diverse methods under a single framework enables direct comparisons among them. This comparative lens allows us to better understand their similarities, differences, and behavior, such as why different methods can generate disagreeing explanations for the same model prediction (Krishna* et al., 2024). Third, this unification enables theoretical simplicity. Instead of studying methods individually, theoretical analyses can be performed using the framework and applied to each method Han et al. (2022); Bilodeau et al. (2024); Bordt and von Luxburg (2023). Fourth, the conceptual understanding brought about by unification leads to principled, practical recommendations (Han et al., 2022). Additional details on this unification are discussed in Appendix C.6.
3.2 Understanding data attributions (DA)
@@
DA studies how the training dataset shapes model behavior during training. It is also known as data valuation, as it assesses the value of data from vendors. For each training data point , an attribution score traces back to training to quantify ’s influence on the model’s output for a test point . DA scores can be used to characterize data properties, help identify mislabeled data, and justify training data values. Like FA, DA uses the techniques of perturbations, gradients, and linear approximation. We examine prominent methods below, with additional details in Appendix D.
@@
Perturbation-based DA. These DA methods examine changes in the model behavior after removing or re-weighting the training data points and subsequently retraining the model (as a result, they are also called retraining-based methods (Hammoudeh and Lowd, 2024)). Leave-One-Out (LOO) attribution, analogous to direct perturbation in FA, is a prominent example that trains a model on the complete training set and then separately removes each individual data point and retrains the model. The attribution score for each removed point is determined by the corresponding performance difference. The LOO approach has a long history in statistics (Cook and Weisberg, 1982) and has proven useful for DA for modern AI models (Jia et al., 2021). It provides valuable insights with its main limitation being the computational cost of model retraining for each data point. Many DA methods later can be viewed as efficient approximations of LOO. A natural extension of LOO is to leave a set of data points out to evaluate their collective impact (Ilyas et al., 2022). As the direct perturbation for FA, LOO overlooks interactions between data points, potentially missing subtle influence behaviors (Lin et al., 2022; Jia et al., 2021). This is addressed by game-theoretic DA methods like Data Shapley (Ghorbani and Zou, 2019) which applies the Shapley value to DA by computing each training point’s aggregated marginal contribution across all training subsets. Similar to game-theoretic FA, they face prohibitive computational costs, as each marginal contribution requires retraining and there are possible subsets. Many approximations have been proposed to address this challenge (Section D.2).
@@
Gradient-based DA. These DA methods use the gradients of the loss with respect to both training data and test data to assess the impact of training point on model output . The GradDot and GradCos methods use the dot product and cosine similarity of these two gradients as attribution scores Charpiat et al. (2019). Influence Function (IF), a classic statistical method for analyzing influential points in linear regression (Cook and Weisberg, 1980), has been adapted for modern AI models (Koh and Liang, 2017). IF approximates LOO model parameter changes by Taylor expansion, computing both the gradient and the (inverse) Hessian, thus avoiding retraining. It offers an effective and computationally feasible alternative to LOO but also faces challenges including convexity assumptions and Hessian matrix computation which many follow-up methods try to address (Section D.3). Aside from these IF-based methods, TracIn (Pruthi et al., 2020) attributes the data influence throughout the entire training path by computing dot products between training and test data gradients at each training step from the initial model parameters to the final model parameters, accumulating these to capture a total influence across the training path. This path tracing method avoids limitations of LOO and IF, such as assigning identical attribution scores to duplicate training data points.
@@
Linear approximation for DA. Datamodel (Ilyas et al., 2022) applies linear approximation to DA, similar to LIME in FA. It constructs a linear model with coefficients and vectors as inputs, where each input represents a subset of training data. is learned to map any subset of training data to output , where is trained on this subset with the given model architecture and training algorithm. The coefficients of thus represent the attribution scores of the data points. While Datamodel can effectively capture model behavior, constructing this large linear model with coefficients requires extensive counterfactual data obtained by training model on various subsets, making it computationally intensive. TRAK (Park et al., 2023) addresses the computational challenge by estimating Datamodels in a transformed space where the learning problem becomes convex. Furthermore, both approaches can be viewed as perturbation-based methods, similar to LIME, as they vary training data to construct linear models.
3.3 Understanding component attributions (CA)